# Strong-to-Weak Spontaneous Symmetry Breaking in Mixed Quantum States

### v1

@@ Appendix A $\mathrm{U}(1)$ Rényi-2 SW-SSB in the rotor model
$\operatorname{tr}\rho^{2}\propto\int_{\theta,\theta^{\prime}}\prod_{\langle ij\rangle}F^{2}_{ij}(\theta,\theta^{\prime}).$ (93)
@@
Appendix B The recovery channel
@@
B.1 Recovery of a local channel
@@ B.3 Global recovery
The strategy of the global recovery is to recover $\mathcal{E}$ layer by layer. We cut the system into hypercubes of linear size $L_{0}$ – thus we have $(L/L_{0})^{d}$ hypercubes in total, where $d$ is the spatial dimension of the system. Here $L_{0}$ is assumed to be larger than $w$ and the correlation length $1/2\Delta$. We first reorganize each layer of $\mathcal{E}$ into $2^{d}$ new layers. In other words, we now perceive the channel $\mathcal{E}$ as a circuit with $h=2^{d}l$ layers, but the gates become “sparse.” Specifically, in each new layer, within a hypercube of size $2L_{0}$, only a hypercube of size $L_{0}$ contains nontrivial gates. Crucially, for each new layer, these nontrivial hypercubes are separated by a minimum distance of $L_{0}$. In the first recovery step, we reverse all gates in the last ($h$-th) layer of the new circuit. We label the hypercubes undergoing recovery as $C_{x}$ with $x\in{1,\dots,(L/2L_{0})^{d}}$, the gates in the $h$-th layer within $C_{x}$ as $\mathcal{E}_{x}^{h}$, and the corresponding recovery channels as ${\mathcal{R}^{h}_{x}}$. The recovery map $\mathcal{R}^{h}_{x}$ is chosen to be the local recovery map mentioned in Sec B.1, confined to a hypercube $C_{x}^{\prime}$ with a linear size of $2L_{0}$, and sharing the same center as the hypercube $x$. Note that these recovery channels act on disjoint hypercubes, and can therefore be implemented simultaneously in a single layer. Once more, we denote the state after each new layer of $\mathcal{E}$ as $\rho_{k}$, now with $0\leq k\leq 2^{d}l$. We have:

### v2

@@ Appendix C $\mathrm{U}(1)$ Rényi-2 SW-SSB in the rotor model
$\operatorname{tr}\rho^{2}\propto\int_{\theta,\theta^{\prime}}\prod_{\langle ij\rangle}G^{2}_{ij}(\theta,\theta^{\prime}).$ (110)
@@
Appendix D The recovery channel
@@
D.1 Recovery of a local channel
@@ D.3 Global recovery
The strategy of the global recovery is to recover $\mathcal{E}$ layer by layer. We cut the system into hypercubes of linear size $L_{0}$ – thus we have $(L/L_{0})^{d}$ hypercubes in total, where $d$ is the spatial dimension of the system. Here $L_{0}$ is assumed to be larger than $w$ and the correlation length $1/2\Delta$. We first reorganize each layer of $\mathcal{E}$ into $2^{d}$ new layers. In other words, we now perceive the channel $\mathcal{E}$ as a circuit with $h=2^{d}l$ layers, but the gates become “sparse.” Specifically, in each new layer, within a hypercube of size $2L_{0}$, only a hypercube of size $L_{0}$ contains nontrivial gates. Crucially, for each new layer, these nontrivial hypercubes are separated by a minimum distance of $L_{0}$. In the first recovery step, we reverse all gates in the last ($h$-th) layer of the new circuit. We label the hypercubes undergoing recovery as $C_{x}$ with $x\in{1,\dots,(L/2L_{0})^{d}}$, the gates in the $h$-th layer within $C_{x}$ as $\mathcal{E}_{x}^{h}$, and the corresponding recovery channels as ${\mathcal{R}^{h}_{x}}$. The recovery map $\mathcal{R}^{h}_{x}$ is chosen to be the local recovery map mentioned in Sec D.1, confined to a hypercube $C_{x}^{\prime}$ with a linear size of $2L_{0}$, and sharing the same center as the hypercube $x$. Note that these recovery channels act on disjoint hypercubes, and can therefore be implemented simultaneously in a single layer. Once more, we denote the state after each new layer of $\mathcal{E}$ as $\rho_{k}$, now with $0\leq k\leq 2^{d}l$. We have: